OpenStudy (anonymous):

Use combinations of Properties 1 and 2 for radicals to simplify the following problem as much as possible. Assume the variables represent positive numbers. @kx2bay

OpenStudy (anonymous):

\[\frac{ 4\sqrt{27x^2y} }{ \sqrt{9} }\]

OpenStudy (anonymous):

did you give it a try cookii?

OpenStudy (anonymous):

did you give it a try cookii?

OpenStudy (anonymous):

\[4xy \sqrt{3}\]

OpenStudy (anonymous):

Wait, no the y stays in the radical

OpenStudy (anonymous):

\[4x \sqrt{3y}\]

OpenStudy (anonymous):

@kx2bay , where'd you go? Are you watching Family Guy? lol

OpenStudy (anonymous):

hahaha hate this f-n guy

OpenStudy (anonymous):

hahaha hate this f-n guy

OpenStudy (anonymous):

lol family guy is awesome lol

OpenStudy (anonymous):

u too smart now cookii what's going on ?

OpenStudy (anonymous):

u too smart now cookii what's going on ?

OpenStudy (anonymous):

lol idk

OpenStudy (anonymous):

i think I know... you decided to think for yourself and just proven that you can get the right answer with a bit of confidence :)

OpenStudy (anonymous):

keep it up cookii, you going places...